# Attention-Aided MMSE for OFDM Channel Estimation: Learning Linear Filters with Attention

### v1

@@ IV-C1 MSE and Huber Loss
To improve robustness to noise and extreme outliers, and better reflect channel statistics, we train A-MMSE with the Huber loss, which stabilize learning. The corresponding methodology is explained in detail in Section VI.
@@ V Rank-Adaptive A-MMSE
Figure 5: 5G NR time-frequency resource block with Type A DM-RS.
@@
[…]
Considering an OFDM channel comprising $6$ RBs (minimum bandwidth assignment), i.e., $N=72$, $M=14$, and $L=36\ \text{or}\ 72$, the RA-A-MMSE with $r=12$ has the reduced computational cost by approximately $35\%$ and $18\%$ for $L=36$ and $L=72$, respectively. Moreover, since the RA module leverages learnable matrices $\left(\mathbf{U}_{r}\ \text{and}\ \mathbf{V}_{r}\right)$ optimized during training, this substantial complexity reduction can be achieved with minimal channel estimation performance loss. We demonstrate this in the next section.
@@ VI Simulation Results
VI-A Simulation Setup
We evaluate channel estimation performance using TDL channel models (TDL-A to TDL-E) specified in 3GPP TR 38.901, which capture realistic multipath propagation scenarios (see Table I). Model training is conducted using TensorFlow on a single NVIDIA GeForce RTX 4090 GPU.
@@ VI-A Simulation Setup
Due to space limitations, we focus on three representative TDL models: TDL-B (typical urban), TDL-D (high mobility), and TDL-E (mmWave), which together span a broad range of deployment scenarios across frequency bands. As benchmark methods, we consider the following. For SP-based approaches, we include LS, MMSE, and 1D-MMSE. For DNN-based approaches, we include ChannelNet [7] and Channelformer [13].
@@
VI-B Normalized MSE Comparison
For the channel estimation performance comparison, we present the NMSE across different SNR levels in Fig. 6 and Fig. 7. As shown in the figures, the proposed A-MMSE method consistently achieves the best NMSE performance across all SNR regimes and propagation environments, clearly outperforming the existing SP-based approaches as well as the existing DNN-based approaches. Specifically, the A-MMSE achieves around $43\%$ and $48\%$ lower average NMSE compared to the best SP-based baseline (MMSE) and the best DNN-based baseline (Channelformer), respectively. In the considered TDL models, channel statistics vary from slot to slot, rendering classical MMSE no longer strictly optimal. In this situation, the A-MMSE still achieves significantly lower NMSE than MMSE, where the gains are more pronounced in the TDL-D and TDL-E (harsh high-mobility and frequency selective conditions). This advantage arises from that the A-MMSE leverages the powerful learning capability from the Attention Transformer. It is also notable that the proposed A-MMSE outperforms Channelformer, which also employs the Attention Transformer architecture. These gains arise from the novel A-MMSE’s two-stage Attention encoder, which first captures frequency-domain correlations and then refines them in the temporal-domain correlation. Crucially, this enables learning the non-stationary channel statistics that vary across OFDM slots more effectively compared to the single-encoder design used in Channelformer.
@@ VI-B Normalized MSE Comparison
VI-C Robustness to SNR Mismatch
To evaluate the robustness and generalization of the proposed A-MMSE under SNR mismatch conditions, we consider three independent models trained at SNR values of 00 dB, $15$ dB, and $30$ dB, and evaluate each model across the entire SNR range of $0-30$ dB. Fig. 8 and Fig. 9(a) illustrate the NMSE performance of A-MMSE across the TDL-B, TDL-D, and TDL-E scenarios, respectively.
Across all considered channel profiles, the A-MMSE with SNR mismatch reveals a consistent trend. For instance, the A-MMSE model trained at 00 dB collapses when the test SNR increases. This is mainly because the overwhelming noise obscures the channel’s latent statistics, causing the Attention Transformer to memorize a denoiser rather than learn the true optimal filter. In contrast, the A-MMSE models trained at $15$ dB and $30$ dB generalize effectively to unseen SNR levels. Specifically, in TDL-B, the $15$ dB model retains $99.8\%$ and the $30$ dB model retains $99.3\%$ of the no-mismatch reference performance. In TDL-D, the corresponding performance retentions are $99.1\%$ and $98.1\%$, and in TDL-E, they are $99.0\%$ and $98.1\%$, respectively.
We find that the persistent advantage of the $15$ dB model stems from a sweet-spot effect, whereby moderate noise serves as regularization during the learning process. This allows the network to effectively internalize the full MMSE filter structure without overfitting to an idealized, noise-free setting. Compared to this, at $30$ dB, the regularization effect from noise becomes nearly negligible, resulting in a slight loss of robustness as operating conditions deviate from the training SNR.
@@ VI-C Robustness to SNR Mismatch
VI-D Performance-Complexity Trade‑off
@@ VII Conclusion
The proposed A-MMSE and the RA-A-MMSE frameworks offer several key advantages: i) once trained, only the linear filter is used, resulting in architectural simplicity; ii) by leveraging the learning capability of the Attention Transformer, it adapts effectively to diverse channel conditions; and iii) it provides a principled integration of classical signal processing with deep learning. With its modular design and computational efficiency, the A-MMSE aligns well with emerging 6G physical layer design principles. Its integration of data-driven learning with model-based filtering makes it well-suited for standardization in scalable, low-complexity, and AI-native wireless transceiver design. Consequently, the proposed architecture offers a practical pathway toward learning-enhanced, standards-compliant channel estimation in next-generation wireless systems.

### v2

@@ IV-C1 MSE and Huber Loss
Figure 4: 5G NR time-frequency resource block with Type A DM-RS.
To improve robustness to noise and extreme outliers, and better reflect channel statistics, we train A-MMSE with the Huber loss, which stabilizes learning. As demonstrated in Fig. 4, Huber loss consistently outperform MSE loss, particularly at high SNR where its bounded gradient prevents occasional large residuals from destabilizing the training process-achieving up to $17\%$ lower NMSE at $30$ dB SNR. The detailed training methodology using Huber loss and its implementation for channel estimation are explained in Section VI.
@@ V Rank-Adaptive A-MMSE
Figure 6: 5G NR time-frequency resource block with Type A DM-RS.
@@
$\text{Complexity Cost Ratio:}\;\frac{\text{RA-A-MMSE}}{\text{A-MMSE}}\approx\frac{NMr+Lr}{NML}$ (36)
Considering an OFDM channel comprising $6$ RBs (minimum bandwidth assignment), i.e., $N=72$, $M=14$, and L=36or 72L=36\ $\text${or}\ 72, the RA-A-MMSE with $r=12$ has the reduced computational cost by approximately $35\%$ and $18\%$ for $L=36$ and $L=72$, respectively. Moreover, since the RA module leverages learnable matrices $\left(\mathbf{U}_{r}\ \text{and}\ \mathbf{V}_{r}\right)$ optimized during training, this substantial complexity reduction can be achieved with minimal channel estimation performance loss. We demonstrate this in the next section.
Remark 4.
The proposed A-MMSE method exploits sparsity in two primary ways. First, the two-stage Attention encoder implicitly leverages sparsity through its MHA mechanism, which naturally induces soft-sparsity by assigning higher weights to dominant features while suppressing less significant ones. This aligns well with wideband wireless channels that can be captured by a separable, low-rank covariance model ($\mathbf{R}_{\mathrm{full}}\approx\mathbf{R}_{f}\otimes\mathbf{R}_{t}$). Second, the Rank-Adaptive A-MMSE (RA-A-MMSE) framework exploits the rank-deficient structure inherent in the ideal MMSE filter, where all rows share a common whitening filter and differ only by position-dependent correlation coefficients. By learning to identify and operate within this naturally sparse subspace, RA-A-MMSE achieves dramatic rank reduction without significant performance loss. As shown in Section VI, RA-A-MMSE maintains $99.6\%$ of its performance with only $10\%$ of the full rank, demonstrating that it identifies the naturally sparse subspace where MMSE channel estimation resides rather than simply discarding information.
@@ VI Simulation Results
Figure 9: BER performance under 16-QAM modulation for the 3GPP TDL-E channel model.
VI-A Simulation Setup
We evaluate channel estimation performance using TDL channel models (TDL-A to TDL-E) specified in 3GPP TR 38.901, which capture realistic multipath propagation scenarios (see Table I). For each scenario, a continuous sequence of $44,000$ OFDM frames was generated using a standard channel model simulator. By running the channel model continuously without resets and dynamically updating the Doppler shift to model user mobility, we intentionally created a non-stationary channel where the WSSUS assumption is broken globally and holds only locally over short intervals. The generated dataset was partitioned to test generalization: the last 4,000 frames served as the test set with temporal separation from the initial $40,000$ frames, which were divided into $36,000$ training frames and $4,000$ validation frames. This temporal split ensures natural statistical drift between training and test sets, providing a realistic measure of robustness. Model training was conducted using TensorFlow on a single NVIDIA GeForce RTX 4090 GPU.
Due to space limitations, we focus on three representative TDL models: TDL-B (typical urban), TDL-D (high mobility), and TDL-E (mmWave), which together span a broad range of deployment scenarios across frequency bands. The diverse subcarrier spacings specified in Table I, corresponding to different symbol durations, are chosen intentionally to evaluate the proposed method across the fundamental trade-off between Doppler resolution and robustness to channel time-variations. As benchmark methods, we consider the following. For SP-based approaches, we include LS, MMSE, and 1D-MMSE. For DNN-based approaches, we include ChannelNet [9] and Channelformer [15]. Furthermore, in the final subsection, we also compare the performance of a sensing-aided approach [7] and A-MMSE.
@@ VI-A Simulation Setup
VI-B On the Suboptimality of Covariance-Mismatched MMSE Estimation
As described in subsection VI-A, our dataset was intentionally generated to break the WSSUS assumption globally, creating a non-stationary environment where channel statistics vary across OFDM frames. This directly causes covariance mismatch in conventional MMSE estimation, as the filter relies on a fixed estimated covariance $\tilde{\mathbf{R}}$ while the true covariance $\mathbf{R}$ changes over time. The conventional MMSE estimator suffers from an unavoidable performance penalty due to this mismatch, quantified as
$\Delta_{\mathrm{mis}}=\left\|(\mathbf{W}_{\mathrm{mis}}-\mathbf{W}^{*})\Sigma_{yy}^{1/2}\right\|_{F}^{2}\geq 0.$ (37)
where $\mathbf{W}_{\mathrm{mis}}$ denotes the mismatched filter and $\mathbf{W}^{*}$ represents the oracle MMSE filter with perfect covariance knowledge. As detailed in Appendix B, this penalty is proportional to $\lVert\Delta\mathbf{R}\rVert^{2}_{F}$ and establishes a fundamental performance floor, particularly severe in high-mobility scenarios such as TDL-D and TDL-E.
In contrast, A-MMSE learns the optimal filter directly from data without explicit covariance estimation, bypassing the mismatch problem entirely. The attention mechanism captures true non-stationary channel correlations through an adaptive filter whose error decreases with model capacity and training samples, unlike the fixed penalty $\Delta_{\text{mis}}$ in conventional MMSE. This fundamental distinction explains A-MMSE’s consistent advantage, achieving approximately $43\%$ lower NMSE compared to conventional MMSE across all TDL scenarios.
VI-C Normalized MSE Comparison
For the channel estimation performance comparison, we present the NMSE across different SNR levels in Fig. 7 and Fig. 8. As shown in the figures, the proposed A-MMSE method consistently achieves the best NMSE performance across all SNR regimes and propagation environments, clearly outperforming the existing SP-based approaches as well as the existing DNN-based approaches. Specifically, the A-MMSE achieves around $43\%$ and $48\%$ lower average NMSE compared to the best SP-based baseline (MMSE) and the best DNN-based baseline (Channelformer), respectively. In the considered TDL models, channel statistics vary from slot to slot, rendering classical MMSE no longer strictly optimal. In this situation, the A-MMSE still achieves significantly lower NMSE than MMSE, where the gains are more pronounced in the TDL-D and TDL-E (high-mobility and mmWave). This advantage arises from that the A-MMSE leverages the powerful learning capability from the Attention Transformer. It is also notable that the proposed A-MMSE outperforms Channelformer, which also employs the Attention Transformer architecture. These gains arise from the novel A-MMSE’s two-stage Attention encoder, which first captures frequency-domain correlations and then refines them in the temporal-domain correlation. Crucially, this enables learning the non-stationary channel statistics that vary across OFDM slots more effectively compared to the single-encoder design used in Channelformer.
@@ VI-C Normalized MSE Comparison
VI-D BER Performance Analysis
To evaluate the impact of channel estimation accuracy on symbol detection performance, we examine the BER performance for 16-QAM modulation in the TDL-E scenario. Fig. 9 presents the BER curves for the challenging mmWave channel environment. The results reveal distinct performance characteristics at different SNR regimes. At low SNRs (0 - $10$ dB), all methods exhibit similar BER performance, as the noise dominates the overall system performance. However, at high SNRs ($20$ - $30$ dB), where channel estimation accuracy becomes the limiting factor, the A-MMSE demonstrates substantial gains. At $25$ dB, the A-MMSE achieves a BER of $8.59\times 10^{-5}$ compared to $9.89\times 10^{-5}$ for MMSE and $1.17\times 10^{-4}$ for Channelformer, representing improvements of $13.1\%$ and $26.6\%$ respectively. These BER results confirm that the superior channel estimation accuracy of A-MMSE directly translates to enhanced symbol detection performance, with benefits becoming increasingly significant in high SNR.
VI-E Robustness to SNR Mismatch
To evaluate the robustness and generalization of the proposed A-MMSE under SNR mismatch conditions, we consider three independent models trained at SNR values of 0 dB, $15$ dB, and $30$ dB, and evaluate each model across the entire SNR range of 0300-30 dB. Fig. 10 and Fig. 11(a) illustrate the NMSE performance of A-MMSE across the TDL-B, TDL-D, and TDL-E scenarios, respectively.
Across all considered channel profiles, the A-MMSE with SNR mismatch reveals a consistent trend. For instance, the A-MMSE model trained at 0 dB collapses when the test SNR increases. This is mainly because the overwhelming noise obscures the channel’s latent statistics, causing the Attention Transformer to memorize a denoiser rather than learn the true optimal filter. In contrast, the A-MMSE models trained at $15$ dB and $30$ dB generalize effectively to unseen SNR levels. Specifically, in TDL-B, the $15$ dB model retains $99.8\%$ and the $30$ dB model retains $99.3\%$ of the no-mismatch reference performance. In TDL-D, the corresponding performance retentions are $99.1\%$ and $98.1\%$, and in TDL-E, they are $99.0\%$ and $98.1\%$, respectively.
The persistent advantage of the ($15$) dB model stems from a sweet-spot effect, where moderate noise serves as regularization, allowing the network to internalize the full MMSE filter structure without overfitting to noise-free conditions. At ($30$) dB, the regularization effect becomes nearly negligible, resulting in slight robustness loss as operating conditions deviate from the training SNR.
VI-F Performance-Complexity Trade‑off
@@
VI-G Performance Comparison to Sensing-aided Estimation
Recent sensing-aided channel estimation methods present significant trade-offs compared to the A-MMSE framework, particularly in high-mobility scenarios like TDL-D. The sensing-aided approach requires a complex multi-stage processing pipeline for echo analysis, adding considerable computational overhead. Unlike this, A-MMSE performs channel estimation through a single, efficient linear operation. Moreover, sensing-aided techniques depend critically on echo signal quality, a vulnerability absent in A-MMSE, which maintains consistent performance without auxiliary information. In Fig. 12, we also compare the channel estimation performance between our A-MMSE and the sensing-aided approach. At $15$ dB SNR in TDL-D, A-MMSE achieves an NMSE of $3.26\times 10^{-3}$, outperforming the sensing-aided approach ($5.01\times 10^{-3}$) by $35\%$. This performance gap widens at higher SNRs, where sensing-aided methods plateau while A-MMSE continues to improve. It demonstrates that data-driven learning overcomes the fundamental limitations of echo-based refinement. These results indicate that A-MMSE retains its advantage even without relying on explicit echo information, highlighting its robustness and applicability.
Figure 12: A-MMSE vs. Sensing-aided channel estimation
@@ VII Conclusion
The proposed A-MMSE and the RA-A-MMSE frameworks offer several key advantages: i) once trained, only the linear filter is used, resulting in architectural simplicity; ii) by leveraging the learning capability of the Attention Transformer, it adapts effectively to diverse channel conditions; and iii) it provides a principled integration of classical signal processing with deep learning. With its modular design and computational efficiency, the A-MMSE aligns well with emerging 6G physical layer design principles. Its integration of data-driven learning with model-based filtering makes it well-suited for standardization in scalable, low-complexity, and AI-native wireless transceiver design. Consequently, the proposed architecture offers a practical pathway toward learning-enhanced, standards-compliant channel estimation in next-generation wireless systems. Future work could explore a promising hybrid model that integrates the A-MMSE framework with sensing-aided approaches, while also developing a practical ground-truth-free training methodology using Ridge regression. These advancements would lay the groundwork for a fully online A-MMSE framework that can adapt in real-time to dynamic wireless environments without perfect channel knowledge.
Appendix A Proof of Kronecker Product Structure
Consider the time-varying channel impulse response $h(t,\tau)$ [39] with time-frequency transfer function
$L_{H}(t,f)=\int^{\infty}_{-\infty}{h(t,\tau)e^{-j2\pi f\tau}\,d\tau}.$ (38)
Under WSSUS [39]: $\mathbb{E}\left[h(t,\tau)h^{*}(t^{\prime},\tau^{\prime})\right]=r_{h}(t-t^{\prime};\tau)\delta(\tau-\tau^{\prime}).$ This yields 2D stationarity:
$\mathbb{E}\!\left[L_{H}(t,f)L^{*}_{H}(t^{\prime},f^{\prime})\right]$ (39)
$=\iint\mathbb{E}\!\left[h(t,\tau)h^{*}(t^{\prime},\tau^{\prime})\right]\,e^{-j2\pi f\tau}e^{j2\pi f^{\prime}\tau^{\prime}}\,d\tau d\tau^{\prime}$ (40)
$=\int r_{h}(t-t^{\prime};\tau)\,e^{-j2\pi(f-f^{\prime})\tau}\,d\tau.$ (41)
Setting $\Delta t=t^{\prime}-t$, $\Delta f=f^{\prime}-f$, then we have $\mathbb{E}\left[L_{H}(t,f)L_{H}^{*}(t+\Delta t,f+\Delta f)\right]=R_{H}(\Delta t,\Delta f)$. This correlation function and scattering function form a 2D Fourier pair:
$R_{H}(\Delta t,\Delta f)$ $=\int^{\infty}_{-\infty}{\int^{\infty}_{-\infty}C_{H}(\tau,\nu){e^{j2\pi(\nu\Delta t-\tau\Delta f)\,d\tau d\nu}}}.$ (42)
Assuming separable scattering [35] $C_{H}(\tau,\nu)=P_{\tau}(\tau)\cdot S_{\nu}(\nu)$, we can substitute as
$R_{H}(\Delta t,\Delta f)$ $=\int^{\infty}_{-\infty}{\int^{\infty}_{-\infty}{P_{\tau}(\tau)S_{\nu}(\nu)e^{j2\pi(\nu\Delta t-\tau\Delta f)}}}\,d\tau d\nu$ (43)
$=\int^{\infty}_{-\infty}{P_{\tau}(\tau)e^{-j2\pi\tau\Delta f}\,d\tau}\cdot\int^{\infty}_{-\infty}{S_{\nu}(\nu)e^{j2\pi\nu\Delta t}\,d\nu}$ (44)
$=r_{t}(\Delta t)\cdot r_{f}(\Delta f).$ (45)
For OFDM with $N$ subcarriers and $M$ symbols: $H[n.m]=L_{H}(mT_{sym},\,n\Delta f_{s})$. The discrete correlation: $\mathbb{E}\left[H[n,m]H^{*}[n^{\prime},m^{\prime}]\right]=r_{t}\left[(m-m^{\prime})T_{sym}\right]\cdot r_{f}[(n-n^{\prime})\Delta f_{s}]$ Define frequency and temporal correlation matrices as
  • $\left[\mathbf{R}_{f}\right]_{n,n^{\prime}}=r_{f}\left[(n-n^{\prime})\Delta f_{s}\right]\in\mathbb{C}^{N\times N}$
  • $\left[\mathbf{R}_{t}\right]_{m,m^{\prime}}=r_{t}\left[(m-m^{\prime})T_{sym}\right]\in\mathbb{C}^{M\times M}$
For vectorized channel $\mathbf{h}=\operatorname{vec}\left(\mathbf{H}\right)$ with $h_{n+mN}=H[n,m]$, OFDM channel’s correlation can be represented
$[\mathbf{R}]_{(n+mN),(n^{\prime}+m^{\prime}N)}$ $=\mathbb{E}\left[H[n,m]H^{*}[n^{\prime},m^{\prime}]\right]$ (46)
$=[\mathbf{R}_{t}]_{m,m^{\prime}}\cdot[\mathbf{R}_{f}]_{n,n^{\prime}}.$ (47)
Therefore $\mathbf{R}=\mathbb{E}\left[\mathbf{h}\mathbf{h}^{H}\right]=\mathbf{R}_{t}\otimes\mathbf{R}_{f}$.
Appendix B Performance Penalty of Covariance-Mismatched MMSE
In this appendix, we provide a rigorous mathematical analysis of the performance penalty incurred by covariance-mismatched MMSE estimation. The oracle MMSE filter, assuming perfect knowledge of the true channel covariance $\mathbf{R}$, is given by
$\mathbf{W}^{*}=\Sigma_{xy}\Sigma_{yy}^{-1}=\mathbf{R}\mathbf{S}^{H}\left(\mathbf{S}\mathbf{R}\mathbf{S}^{H}+\sigma^{2}\mathbf{I}\right)^{-1}.$ (48)
where $\mathbf{S}$ denotes the pilot selection matrix, $\Sigma_{xy}=\mathbf{R}\mathbf{S}^{H}$ is the cross-covariance between the full channel and pilot observations, and $\mathbf{S}\mathbf{R}\mathbf{S}^{H}+\sigma^{2}\mathbf{I}$ is the auto-covariance of pilot observations. The corresponding minimum achievable MSE is
$\mathrm{MSE}(\mathbf{W}^{\star})$ $=\mathrm{tr}(\mathbf{R})-\mathrm{tr}\big(\Sigma_{xy}\Sigma_{yy}^{-1}\Sigma_{yx}\big).$ (49)
In practice, we must rely on an estimated or surrogate covariance $\tilde{\mathbf{R}}$, leading to the mismatched filter
$\mathbf{W}_{\mathrm{mis}}=\tilde{\mathbf{R}}\mathbf{S}^{H}\left(\mathbf{S}\tilde{\mathbf{R}}\mathbf{S}^{H}+\sigma^{2}\mathbf{I}\right)^{-1}.$ (50)
The performance of this mismatched filter is strictly worse than the oracle, with total MSE given by
$\mathrm{MSE}(\mathbf{W}_{\mathrm{mis}})=\mathrm{MSE}(\mathbf{W}^{*})+\Delta_{\mathrm{mis}}.$ (51)
where $\Delta_{\mathrm{mis}}\geq 0$ represents the irreducible penalty due to covariance mismatch. The penalty term can be explicitly quantified as
$\Delta_{\mathrm{mis}}=\left\|(\mathbf{W}_{\mathrm{mis}}-\mathbf{W}^{*})\Sigma_{yy}^{1/2}\right\|_{F}^{2}\geq 0.$ (52)
To prove this, we expand the MSE of the mismatched filter:
$\mathrm{MSE}(\mathbf{W}_{\mathrm{mis}})=\mathbb{E}\!\left[\left\|\mathbf{H}-\mathbf{W}_{\mathrm{mis}}\mathbf{y}_{p}\right\|_{2}^{2}\right].$ (53)
where $\mathbf{H}$ is the true channel vector and $\mathbf{y}_{p}$ is the pilot observation vector. By adding and subtracting $\mathbf{W}^{\star}\mathbf{y}_{p}$, we obtain
$\mathrm{MSE}(\mathbf{W}_{\mathrm{mis}})=\mathbb{E}\!\left[\left\|(\mathbf{H}-\mathbf{W}^{*}\mathbf{y}_{p})+(\mathbf{W}^{*}-\mathbf{W}_{\mathrm{mis}})\mathbf{y}_{p}\right\|_{2}^{2}\right].$ (54)
Using the orthogonality principle of the optimal MMSE estimator, the cross term vanishes, yielding
$\mathrm{MSE}(\mathbf{W}_{\mathrm{mis}})=\mathrm{MSE}(\mathbf{W}^{*})+\mathbb{E}\!\left[\left\|(\mathbf{W}^{*}-\mathbf{W}_{\mathrm{mis}})\mathbf{y}_{p}\right\|_{2}^{2}\right].$ (55)
The second term can be written as
$\Delta_{\mathrm{mis}}$ $=\operatorname{tr}\!\left((\mathbf{W}_{\mathrm{mis}}-\mathbf{W}^{*})\boldsymbol{\Sigma}_{yy}(\mathbf{W}_{\mathrm{mis}}-\mathbf{W}^{*})^{\sf{H}}\right).$ (56)
which is equivalent to $\Delta_{\mathrm{mis}}=\left\|(\mathbf{W}_{\mathrm{mis}}-\mathbf{W}^{*})\Sigma_{yy}^{1/2}\right\|_{F}^{2}\geq 0.$
For small covariance mismatches $\Delta\mathbf{R}=\tilde{\mathbf{R}}-\mathbf{R}$, we can establish a local quadratic lower bound. Let $\tilde{{\Sigma}}_{yy}=\mathbf{S}\tilde{\mathbf{R}}\mathbf{S}^{\sf{H}}+\sigma^{2}\mathbf{I}.$ and $\Delta{\Sigma}_{yy}=\tilde{{\Sigma}}_{yy}-{\Sigma}_{yy}=S\,\Delta\mathbf{R}\,\mathbf{S}^{\sf{H}}.$. Using first-order perturbation analysis of the matrix inverse, we have
$\mathbf{W}_{\text{mis}}-\mathbf{W}^{*}\approx\Delta\mathbf{R}\mathbf{S}^{\sf{H}}{\Sigma}_{yy}^{-1}-\mathbf{R}\mathbf{S}^{\sf{H}}{\Sigma}_{yy}^{-1}\Delta{\Sigma}_{yy}{\Sigma}_{yy}^{-1}.$ (57)
Substituting $\Delta{\Sigma}_{yy}=\mathbf{S}\,\Delta\mathbf{R}\,\mathbf{S}^{\sf{H}}$ and simplifying as
$\mathbf{W}_{\text{mis}}-\mathbf{W}^{*}\approx(\mathbf{I}-\mathbf{W}^{*}{\mathbf{S}})\,\Delta\mathbf{R}\,\mathbf{S}^{\sf{H}}{\Sigma}_{yy}^{-1}.$ (58)
Therefore, the mismatch penalty satisfies
$\Delta_{\text{mis}}\geq\lambda_{\min}({\Sigma}_{yy})\left\|(\mathbf{I}-\mathbf{W}^{*}\mathbf{S})\,\Delta\mathbf{R}\,\mathbf{S}^{\sf{H}}{\Sigma}_{yy}^{-1/2}\right\|_{F}^{2}.$ (59)
Using the matrix norm property $\lVert\mathbf{AB}\rVert_{F}\geq\sigma_{\mathrm{min}}(\mathbf{A})\lVert\mathbf{B}\rVert_{F}$, where $\sigma_{\mathrm{min}}$ denotes the smallest singular value of a matrix, we obtain
$\Delta_{\text{mis}}\geq\lambda_{\min}({\Sigma}_{yy})\,\sigma_{\min}^{2}(\mathbf{I}-\mathbf{W}^{*}\mathbf{S})\,\sigma_{\min}^{2}(\mathbf{S})\left\|{\Sigma}_{yy}^{-1/2}\Delta\mathbf{R}\right\|_{F}^{2}.$ (60)
Further, noting that $\left\|{\Sigma}_{yy}^{-1/2}\Delta\mathbf{R}\right\|_{F}^{2}\geq\frac{\left\|\Delta\mathbf{R}\right\|_{F}^{2}}{\rho_{\mathbf{S}}^{2}\left\|\mathbf{R}\right\|_{2}+\sigma^{2}}$ for $\mathbf{S}$ with a bounded spectral norm $\rho_{s}=\lVert\mathbf{S}\rVert_{2}$, we establish the quadratic lower bound
$\Delta_{\text{mis}}\geq\lambda_{\min}({\Sigma}_{yy})\,c^{2}\left\|\Delta\mathbf{R}\right\|_{F}^{2},$ (61)
where $c=\frac{\sigma_{\min}(\mathbf{I}-\mathbf{W}^{*}\mathbf{S})\,\sigma_{\min}(\mathbf{S})}{\rho_{s}^{2}\left\|\mathbf{R}\right\|_{2}+\sigma^{2}}>0$ is a positive constant dependent on the system parameters. Even small covariance estimation errors result in an irreducible quadratic performance penalty for conventional signal processing methods.